OpenStudy (anonymous):

Simplify the following: (9x + 7)(3x - 1)

OpenStudy (anonymous):

do you no what FOIL is?

OpenStudy (anonymous):

you should expalian it to them

OpenStudy (anonymous):

true my bad. okay.. (9x+7)(3x-1) you multiply 9x(3x)=27x^2 then 9x(-1)=-9x then 7(3x)= 21x then 7(-1)= -7 so your equation would be 27x^2-9x+21x-7 so if you simplify -9x+21x= 12x so your new equation is 27x^2+12x-7 (sorry I didn't explain it at first)